OpenStudy (anonymous):

how would I write V = x^3 - x^2 - 20x in standard form?

OpenStudy (anonymous):

Do you mean to factor it?

OpenStudy (anonymous):

V=x^3 - x^2 - 20x=x(x^2 -x -20)=x(x-?)(x+?)

OpenStudy (anonymous):

V=x(x-5 ) (4 + x)